OpenStudy (anonymous):

mc088-1.jpg STVU is an isosceles trapezoid. If SV = 3x - 11 and TU = x + 13, find the value of x

OpenStudy (anonymous):

3x-11= x+13 3x-x 13+11 2x=24 x=12